SPECTROSCOPY OF THE I2 MOLECULE IN CRYOGENIC MATRICES OF FLUORINE CONTAINING MOLECULES.

\underline{VADIM A. ALEKSEEV} ^{a}, Insititute of Physics, St.Petersburg State University, 198504 St.Petersburg, Russia; CHRISTINE PRENDERGAST, JOHN G. McCAFFREY, Department of Chemistry, National University of Ireland, Maynooth, Ireland.

The spectroscopy of the I_{2} molecule in CF_{4}, C_{2}F_{6}, C_{3}F_{8} and C_{4}F_{10} cryogenic solids was studied for the first time with use of steady-state one photon excitation in the Vac UV region and two photon laser excitation. In this contribution we discuss spectra and temporal behaviour of emission originating from the ion-pair states of I_{2}. Preliminary results for some other molecular matrices will be presented as well.

aSupport by Science Foundation of Ireland via Walton Visitor Award (07/W.1/I1819) is gratefully acknowledged